OpenStudy (anonymous):

Solve: 6x2 – 15x = 0 a) x= 0 or x= 2.5 b) x= 2 or x= 5 c) x= 2 or x = 5.5 d) x= 0 or x= 2

OpenStudy (mertsj):

Factor out 3x first

OpenStudy (diyadiya):

\[6x^2-15x=0\] 3x is common so factor out 3x What do you get now?

OpenStudy (anonymous):

9

OpenStudy (diyadiya):

9 ? hm ok can you take 3x common from \(6x^2\) and \(15x\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (diyadiya):

Where do you get 2/5=0 from ?

OpenStudy (anonymous):

idk

OpenStudy (diyadiya):

when we take 3x common we get 3x(2x-5)=0

OpenStudy (anonymous):

the answer is a because when you take 3x common.it will be x(2x-5)=0. Thus,x=0 or 2x-5

OpenStudy (diyadiya):

Now 3x=0 and 2x-5=0 can you find x now?

OpenStudy (diyadiya):

@tana can you ?

OpenStudy (anonymous):

yes